# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ VIII Conclusion
In this paper, we developed a hybrid model-based and data-driven method for MOT. The key idea is to use neural networks to enhance specific aspects of the statistical model in model-based Bayesian MOT that have been identified as overly simplistic. The introduced neural-enhanced motion model captures the influence of historical object states and object interactions while the neural-enhanced measurement model makes use of complicated object shape information for data association and false positive measurement rejection. To obtain a good tradeoff between runtime and estimation accuracy, we use sigma points in the neural-enhanced prediction step to evaluate the proposed learnable non-Markovian state transition model efficiently and particles for accurate neural-enhanced measurement update. Our numerical analysis demonstrates that the proposed method efficiently enhances the statistical model used for MOT and, applied to the publicly available nuScenes dataset, achieves state-of-the-art MOT performance. We believe that the proposed hybrid framework can serve as a guideline for the design of future MOT architectures. A potential direction for future research includes an application to underwater tracking problems.

### v3

@@ VIII Conclusion
In this paper, we developed a hybrid model-based and data-driven method for multiobject tracking (MOT.) The key idea is to use neural networks to enhance specific aspects of the statistical model in model-based Bayesian MOT that have been identified as overly simplistic. The introduced neural-enhanced motion model captures the influence of historical object states and object interactions while the neural-enhanced measurement model makes use of complicated object shape information for data association and false positive measurement rejection. To obtain a good tradeoff between runtime and estimation accuracy, we use sigma points (SPs) in the neural-enhanced prediction step to evaluate the proposed learnable non-Markovian state motion model with object interactions efficiently, and particles for accurate neural-enhanced measurement update. Our numerical analysis demonstrates that the proposed method efficiently enhances the statistical model used for MOT and, applied to the publicly available nuScenes dataset, achieves state-of-the-art MOT performance. We believe that the proposed hybrid framework can serve as a guideline for the design of future MOT architectures. A potential direction for future research includes an application to underwater tracking problems.
@@ Acknowledgment
[Uncaptioned image] Shaoxiu Wei (Student Member, IEEE) received the B.E. degree in communication engineering from the University of Electronic Science and Technology of China, Chengdu, China, in 2022. He is currently working toward the Ph.D. degree in statistical signal processing with the University of California San Diego, La Jolla, CA, USA. His research interests include machine learning, statistical signal processing, and wireless sensing.
[Uncaptioned image] Mingchao Liang (Student Member, IEEE) received the B.Eng. degree in electronic engineering from the Chinese University of Hong Kong, Hong Kong SAR, China, in 2018 and the Ph.D. degree in electrical and computer engineering from the University of California San Diego, La Jolla, CA, USA, in 2025. He is currently a Research Scientist at Meta Platforms in Menlo Park, CA, USA. Mr. Liang’s research interests include localization and tracking algorithms, graphical models, machine learning, and deep learning. He was the recipient of a Qualcomm Innovation Fellowship in 2022. Mingchao serves as a reviewer for the IEEE Transactions on Aerospace and Electronic Systems, the IEEE Transactions on Wireless Communications, and the IEEE Transactions on Signal Processing.
[Uncaptioned image] Florian Meyer (Member, IEEE) received the M.Sc. and Ph.D. degrees (with highest honors) in electrical engineering from TU Wien, Vienna, Austria, in 2011 and 2015, respectively. He is an Associate Professor with the University of California San Diego, La Jolla, CA, USA, jointly appointed between the Scripps Institution of Oceanography and the Electrical and Computer Engineering Department. From 2017 to 2019, he was a Postdoctoral Fellow and Associate with the Laboratory for Information & Decision Systems, Massachusetts Institute of Technology, Cambridge, MA, USA, and from 2016 to 2017, he was a Research Scientist with the NATO Centre for Maritime Research and Experimentation, La Spezia, Italy. His research interests include statistical signal processing, high-dimensional and nonlinear estimation, inference on graphs, inverse problems, multiobject tracking (MOT), and simultaneous localization and mapping (SLAM). He is the recipient of the 2021 ISIF Young Investigator Award, a 2022 NSF CAREER Award, a 2022 DARPA Young Faculty Award, and a 2023 ONR Young Investigator Award. He is currently an Associate Editor for IEEE Transactions on Signal Processing. Dr. Meyer also served as an Associate Editor for the IEEE Transactions on Aerospace and Electronic Systems from 2021 to 2023 and the ISIF Journal of Advances in Information Fusion from 2019 to 2022.